Determine the measures of and to the nearest tenth of a degree, and the length of side

AB to one decimal place.


A B

C

3


4

D


A= C=

Side

AB=


1
Expert's answer
2021-06-01T17:45:57-0400

Figure:




here,

"\cos62\degree=\dfrac{l(AB)}{12}\\\ \\\Rightarrow l(AB)=12\times \cos62\degree\\\Rightarrow AB=12\times 0.469 =5.63\ cm"


Hence, "\boxed{AB=6.6\ cm}"